# PAINT: Parallel-in-time Neural Twins for Dynamical System Reconstruction

### v1

@@ 3.3 Theoretical Considerations
Advice for practitioners The selection of an on-trajectory neural twin approach depends on two key factors: (1) Measurement informativeness. Highly informative measurements often justify the use of autoregressive models (e.g., Kalman filters). If measurements lack sufficient information for state reconstruction, parallel-in-time models are preferable to ensure staying on-trajectory – but they require the selected window’s measurements to be informative. (2) System dynamics. If measurements provide limited information, chaotic systems typically demand parallel-in-time models to stay on-trajectory. For stable or periodic systems (e.g., steady-state behavior) autoregressive models may suffice.
@@ 4 Experiments
4.1 Models
PAINT via Flow Matching.
We implement FlowPAINT as a concrete instantiation of PAINT using Flow Matching [2, 3, 46, 49]. PAINT is agnostic to the neural network architecture. We therefore aimed for simplicity and use established building blocks of the computer vision literature [19, 28, 29, 58]. Further experimental details can be found in Appendix A.2.
Data-coupled stochastic interpolants.
PAINT is also agnostic to the generative modeling paradigm, as long as it maps a source to a target distribution. We follow the modeling from Albergo et al. [4] and model the unmasking as data-coupled distribution matching. However, the masking ratio of our method is very high (25 of 16K pixels are unmasked). In early experiments we observed that the gradient signal was too weak for the vicinity of the probes. Therefore, we introduce a spatially weighted loss that puts higher weight on pixels in the neighborhood of the probe points. As stochastic interpolants and Flow Matching in principle lead to the same training setup, we use a common Flow Matching implementation [47].
Autoregressive UNet.
To compare PAINT to an autoregressive model, we use the UNet architecture from Kohl et al. [38]. We follow their proposed paradigm and condition on the probe values $m_{t+1}$ by concatenating the mask and probe values to the channels. The model sizes are chosen to match parameters. FlowPAINT has 19.8M and the autoregressive UNet 20.6M. Further architectural details and hyperparameters are reported in Appendix A.2.
@@ 4.2 Dataset and evaluation
We generate a CFD dataset using the incompressible, pressure-based solver pimpleFOAM in OpenFOAM 8 [54], with subgrid scales modeled by the Smagorinsky LES model [59, 21]. Simulations are performed on a structured mesh, and for training purposes, a subregion is extracted and interpolated onto a regular grid. Since the state-space represents 2D velocities, we will write $\bm{u}_{t}$ instead of $\bm{x}_{t}$ with $\bm{u}_{t}\in\mathbb{R}^{d_{1}\times d_{2}}$. The train/val/test splits across Reynolds numbers are provided in Appendix A.1.
@@
Due to the chaotic and rapidly diverging nature of turbulent flows, comparing a single predicted frame or a short sequence to the ground truth is not meaningful. Research in turbulence instead focuses on the flow’s inherently reproducible statistical properties [59, 18]. Given a possibly infinitely long trajectory $\bm{u}_{[0,t]}$, we follow a common notation in fluid dynamics and decompose $\bm{u}_{[0,t]}$ into a time-averaged mean component $\overline{\bm{u}}\coloneqq\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{i=1}^{t}\bm{u}_{i}$ and a turbulent fluctuation component $\bm{u}^{\prime}_{[0,t]}$, such that $\bm{u}_{[0,t]}=\overline{\bm{u}}+\bm{u}^{\prime}_{[0,t]}$. The variance over time is denoted as $\overline{\bm{u}^{\prime 2}}\coloneqq\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{i=1}^{t}(\bm{u}_{i}-\overline{\bm{u}})^{2}$. Further, we define $E(k)$ as the mean kinetic energy spectrum. While these metrics are defined over an infinite-time, in practice they are computed over a large, finite horizon.
@@
The model was trained with 25 randomly sampled probe points for each data point in a batch. At inference, we investigate the two probe point constellations depicted in Figure 3. Grid: A 10×10 grid of 100 probe points evenly spaced across the domain (see Figure 3(a)). Vertical: A linear arrangement of 25 probe points positioned at the 3/4 axial location of the domain (see Figures 3(b)). In both configurations, an additional probe point is placed at the center of the inlet of the jet to capture the inlet velocity.
@@ 4.3 Results
Model ReRe Probes ${\text{MAE}}(\overline{\mathbf{u}})\downarrow$ ${\text{MAE}}(\overline{\mathbf{u}^{\prime 2}})\downarrow$ $\text{MSE}([\mathbf{\mathbf{u_{[0,t]}}}])\downarrow$ $\text{RMSE}(E(k))\downarrow$
@@
FlowPAINT 2100 grid $1.2\pm 1.6$ $9.9\pm 21.7$ $8.1\pm 3.6$ $6.85$
AR UNet 2100 grid $9.5\pm 14.3$ $98.7\pm 110.4$ $158.3\pm 46.8$ $70.9$
FlowPAINT 2100 vertical $2.2\pm 3.1$ $30.2\pm 38.4$ $49\pm 21$ $20.1$
AR UNet 2100 vertical $9.2\pm 13.8$ $97.8\pm 108.8$ $157.3\pm 70.0$ $71.8$
FlowPAINT 1100 grid $0.9\pm 1.3$ $2.2\pm 5.5$ $2\pm 1$ $1.98$
AR UNet 1100 grid $3.5\pm 5.6$ $17.3\pm 25.9$ $28.8\pm 8.3$ $12.4$
FlowPAINT 1100 vertical $1.1\pm 1.6$ $6.3\pm 11.1$ $8.9\pm 3.4$ $3.68$
AR UNet 1100 vertical $3.4\pm 5.5$ $16.9\pm 25.5$ $28.8\pm 8.1$ $12.2$
Figure 4: Ground truth vs. predicted states at different timesteps for $Re=2100$ using the vertical probe constellation. PAINT (ours) does not accumulate errors over time. For the autoregressive UNet a pronounced drift over time is observable. All values are normalized by the mean inlet velocity.
Figure 5: Mean velocity $\overline{u}$ and variance $\overline{u^{\prime 2}}$ of the reconstructed $Re=2100$ test trajectory using the vertical probe constellation compared with the ground truth. All values are normalized by the mean inlet velocity.
Quantitative results.
The quantitative results are reported in Table 1. FlowPAINT outperforms the autoregressive UNet across all metrics, probe constellations and Reynolds numbers in the test set. We conclude that FlowPAINT is more physically coherent than the probe-conditioned autoregressive baseline. Further results and analyses can be found in Appendix C.
@@
We observe the models ability to reconstruct physical plausible states from different constellations of probe points (see Figures 5 and 6). The model’s ability to reproduce the correct physical statistics seems to depend directly on the number and position of the sample points (as can be seen when comparing Figure 5 with Figures 10(a), 11(a) and 12(a) ).
(a) Time-averaged kinetic energy spectrum.
(b) Histogram of the kinetic energy.
Figure 6: Statistical analysis of reconstructed flow fields for Reynolds number 2100 and a vertical probe point constellation.
(b) Grid probe constellation.
(a) Vertical probe constellation.
Figure 7: MSE over time of ground-truth vs. predicted trajectories for the $Re=2100$ test trajectory and different probe point constellations. Interestingly, the MSE’s of the two autoregressive trajectories are very similar, even if more informative probes are provided for the grid. The autoregressive model indicates over-reliance on the autoregressive state.
Parallel-in-time model maintains stable error, while autoregressive baseline drifts.
The autoregressive approach exhibits a drift in MSE when rolled out (compare Figures 4, 7, 13,14 and 15). Notably, Figure 4 depicts a degradation of the autoregressive model’s state over time, which has been analyzed in prior work [38]. Importantly, the autoregressive model benefits from a known, correct initial state, which is often unknown in practice.
Sampling a sequence vs. a single state.
PAINT can be used in two variants. The first samples a connected sequence, usually $\bm{x}_{[t-h,t]}\sim p(\bm{x}_{[t-h,t])}\mid\bm{m}_{[t-h,t]}$ (see Figure 16). This sequence is smooth over time and can be used by practitioners to interpret results.
The second variant samples a single timestep. It is used for state estimation in a sliding window fashion by sampling $\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{m}_{[t-h,t]})$. In this case a sampled sequence of consecutive states is not smooth over time.
@@
FlowPAINT needs substantially more training resources and is slower at inference than the autoregressive baseline. It was trained on 12 A100 (64 GB VRAM) for 30 hours, whereas the UNet was trained on a single H100 (80GB VRAM) for 24 hours. Notably, the UNet takes about $66$ milliseconds and FlowPAINT around $6.6$ seconds for 20 denoising steps. Further details can be found in Appendix A.2.
@@ 5 Conclusion
In this paper we introduced parallel-in-time neural twins (PAINT) a novel data-driven method for reconstructing dynamical systems from real-time measurements. PAINT employs a generative neural network to model the joint conditional probability of system states. Our method provable stays on-trajectory and does not rely on an initial condition. The empirical validation with FlowPAINT on 2D turbulent fluid dynamics confirms its effectiveness. This paper represents an advancement in designing interpretable and widely applicable neural twins for complex dynamical systems.
@@
The biggest limitation of our method are the high computational costs, which deserves further investigation. Another limitation is that the sliding window approach does not generate continuous generations. This could potentially be mitigated by stitching techniques [68].
@@ Acknowledgements
Appendix A Experimental details
A.1 Dataset
@@
The dataset is generated using the incompressible, pressure-based solver pimpleFOAM in OpenFOAM 8 ([54]), with subgrid scales modeled by the Smagorinsky LES model [59, 21]. Simulations are performed on a structured mesh consisting of 19040 cells, and for training purposes, a subregion is extracted and interpolated onto a regular 128 × 128 grid to facilitate adaptability.. The numerical simulations made use of a timestep size of $\Delta t=6.5\times 10^{-4}\text{s}$. For training purposes, the data are sampled every 70th timestep, resulting in an effective temporal resolution of $4.55\times 10^{-2}\text{s}$. For each Reynolds number, the dataset contains 1000 such snapshots.
@@
In reality, there is no perfectly two-dimensional turbulence; even quasi-2D flows, such as atmospheric flows, exhibit three-dimensional structures at smaller scales. Nevertheless, we are confident that the proposed method, having demonstrated robust performance in two-dimensional turbulence, will generalize effectively to three-dimensional turbulent flows. For a more detailed discussion of two-dimensional turbulence, see [18].

### v2

@@ 4.1 Dataset and evaluation
We generate a CFD dataset using the incompressible, pressure-based solver pimpleFOAM in OpenFOAM 8 (OpenCFD, 2009), with subgrid scales modeled by the Smagorinsky LES model (Pope, 2001; Fröhlich, 2006). Simulations are performed on a structured mesh, and for training purposes, a subregion is extracted and interpolated onto a regular 128 x 128 grid. Since the state-space represents 2D velocities, we will write $\bm{u}_{t}$ instead of $\bm{x}_{t}$ with $\bm{u}_{t}\in\mathbb{R}^{d_{1}\times d_{2}}$. Appendix B.1 provides the train/val/test splits across Reynolds numbers as well as further details about the dataset.
@@
Due to the chaotic and rapidly diverging nature of turbulent flows, comparing a single predicted frame or a short sequence to the ground truth is not meaningful. Research in turbulence instead focuses on the flow’s inherently reproducible statistical properties (Pope, 2001; Davidson, 2015). Given a possibly infinitely long trajectory $\bm{u}_{[0,t]}$, we follow a common notation in fluid dynamics and decompose $\bm{u}_{[0,t]}$ into a time-averaged mean component $\overline{\bm{u}}\coloneqq\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{i=1}^{t}\bm{u}_{i}$ and a turbulent fluctuation component $\bm{u}^{\prime}_{[0,t]}$, such that $\bm{u}_{[0,t]}=\overline{\bm{u}}+\bm{u}^{\prime}_{[0,t]}$. The variance over time is denoted as $\overline{\bm{u}^{\prime 2}}\coloneqq\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{i=1}^{t}(\bm{u}_{i}-\overline{\bm{u}})^{2}$. Further, we define $E(k)$ as the mean kinetic energy spectrum. While these metrics are defined over an infinite-time, in practice they are computed over a large, finite horizon.
@@
The model is trained using 25 random probe points that are resampled for each data point and every iteration. At inference, we investigate the two fixed probe point constellations depicted in Figure 3. Grid: A 10×10 grid of 100 probe points evenly spaced across the domain (see Figure 3(a)). Vertical: A linear arrangement of 25 probe points positioned at the 3/4 axial location of the domain (see Figures 3(b)). In both configurations, an additional probe point is placed at the center of the inlet of the jet to capture the inlet velocity.
Figure 4: Ground truth vs. predicted states at different timesteps for $Re=2100$ using the vertical probe constellation. The prediction labeled PAINT corresponds to FlowPAINT 16/8. Unlike the autoregressive U-Net, PAINT does not accumulate errors over time, whereas the U-Net exhibits a pronounced temporal drift. All values are normalized by the mean inlet velocity.
4.2 Models
PAINT via data-coupled Flow Matching.
PAINT is agnostic to the generative modeling paradigm, as long as it allows for conditional generation. We therefore implement FlowPAINT as a concrete instantiation of PAINT using a common Flow Matching implementation (Lipman et al., 2024). Additionally, the unmasking of non-measurement points in the domain is modeled as data-coupled distribution matching (Albergo et al., 2023b). Notably, the masking ratio of our method is very high (25 of 16K pixels are unmasked). In early experiments we observed that the gradient signal was too weak for the vicinity of the probes. Therefore, we introduce spatial weighting to the Flow Matching loss, which puts higher weight on pixels in the neighborhood of the probe points.
PAINT is also agnostic to the neural network architecture, we therefore use a Diffusion Transformer (DiT) (Peebles & Xie, 2023) with convolutional layers for patching and unpatching. Further experimental details can be found in Appendix B.2.
Autoregressive Baselines.
PAINT is fundamentally different to autoregressive models, because it predicts a sequence of consecutive timesteps at once. In contrast, an autoregressive model predicts the sequence by regressing on its own previous output. Therefore the comparison to an autoregressive model, is inherently difficult. In this work we tested 2 baselines. First, we compare PAINT to the same DiT architecture. However, this setting did not lead to stable rollouts (for further details see Appendix B.2).
To provide a more insightful baseline, we compare PAINT to an established autoregressive UNet baseline from Kohl et al. (2023). The UNet uses convolutional and linear attention (Shen et al., 2021) blocks. We follow their proposed paradigm and condition on two past autoregressive states and on the probe values by concatenating the mask and probe values to the channels. The model sizes were selected to closely match the number of parameters, with FlowPAINT comprising 19.8M parameters and the autoregressive UNet 20.6M. To promote a fair comparison, the optimization hyperparameters and batch sizes were carefully chosen. These and other architectural decisions are explained in further detail in Appendix B.2.
Probe conditioning and future prediction.
Across the experimented models, different measurement histories $h$ and future predictions $n$ are possible (also see Equations 2 and 4). We add the postfix $h+1/n$ to distinguish between these hyperparameters. E.g. “FlowPAINT 16/8” learns the probability $p(\bm{x}_{[t-15,t+8]}\mid\bm{m}_{[t-15,t]})$ of 16 measurement timesteps and predicting 8 timesteps steps into the future. Similarly for autoregressive models, “Autoregressive 16/0” denotes the UNet architecture from Kohl et al. (2023) which learns $p(\bm{x}_{t}\mid\bm{x}_{[t-2,t-1]},\bm{m}_{[t-16,t]})$. Our main results are reported in the 16/0 setting for the best comparability. More settings are reported in the ablations setting in Appendix B.3 and Table 5.
@@ 4.3 Results
Model Cond. ReRe Probes ${\text{MAE}}(\overline{\mathbf{u}})\downarrow$ ${\text{MAE}}(\overline{\mathbf{u}^{\prime 2}})\downarrow$ $\text{MSE}([\mathbf{\mathbf{u_{[0,t]}}}])\downarrow$ $\text{RMSE}(E(k))\downarrow$
@@
FlowPAINT 16/8 2100 grid $1.2\pm 1.6$ $9.9\pm 21.7$ $8.1\pm 3.6$ $6.9$
16/0 $1.3\pm 2.6$ $8.7\pm 15.1$ $10.8\pm 3.9$ $6.3$
Autoreg. 16/0 $8.2\pm 13.7$ $93.0\pm 103.2$ $155.0\pm 50.1$ $66.3$
FlowPAINT 16/8 2100 vertical $2.2\pm 3.1$ $30.2\pm 38.4$ $49.0\pm 21.0$ $20.1$
16/0 $3.0\pm 3.2$ $30.0\pm 41.6$ $68.2\pm 25.3$ $20.5$
Autoreg. 16/0 $7.4\pm 12.5$ $87.6\pm 96.1$ $156.2\pm 50.9$ $62.4$
FlowPAINT 16/8 1100 grid $0.9\pm 1.3$ $2.2\pm 5.5$ $\pm 1.0 2.02.0$
16/0 $1.2\pm 2.8$ $6.2\pm 10.7$ $\pm 1.6 4.64.6$
Autoreg. 16/0 $3.7\pm 6.3$ $16.0\pm 24.1$ $29.9\pm 8.8$ $12.3$
FlowPAINT 16/8 1100 vertical $1.1\pm 1.6$ $6.3\pm 11.1$ $8.9\pm 3.4$ $3.7$
16/0 $3.0\pm 5.8$ $8.2\pm 15.8$ $24.2\pm 6.5$ $9.9$
Autoreg. 16/0 $3.7\pm 6.1$ $15.9\pm 23.9$ $29.5\pm 8.8$ $12.3$
Figure 5: Mean velocity $\overline{u}$ (top row) and variance $\overline{u^{\prime 2}}$ (bottom row) of the reconstructed $Re=2100$ test trajectory using FlowPAINT 16/8 and the vertical probe constellation compared with the ground truth. All values are normalized by the mean inlet velocity.
The quantitative results are reported in Table 1. FlowPAINT outperforms the autoregressive UNet across all metrics, probe constellations and Reynolds numbers in the test set. Further results and ablations can be found in Appendix B.3 and Table 5.
@@
We observe the models ability to reconstruct physical plausible states from different constellations of probe points (see Figures 5 and 16). The model’s ability to reproduce the correct physical statistics seems to depend directly on the number and position of the sample points (as can be seen when comparing Figure 5 with Figures 13(a), 14(a) and 15(a)).
(a) Grid probe constellation.
(b) Vertical probe constellation.
Figure 6: MSE over time of ground-truth vs. predicted trajectories for the $Re=2100$ test trajectory and different probe point constellations, comparing the autoregressive baselines with FlowPAINT. Notably, the autoregressive approach yields nearly identical error profiles for both probe constellations, failing to leverage the additional information provided by the more dense probe placements. The autoregressive model indicates over-reliance on the autoregressive state.
FlowPAINT models maintain stable error, while autoregressive UNet drifts.
The autoregressive approach exhibits a drift in MSE when rolled out (compare Figures 4, 6 and 11). Notably, Figure 4 depicts a degradation of the autoregressive model’s state over time, which has been analyzed in prior work (Kohl et al., 2023). Importantly, the autoregressive model benefits from a known, correct initial state, which is often unknown in practice.
Over-reliance on autoregressive state.
By analyzing the gradients of the loss with respect to the model input, one can estimate how predictions depend on different conditioning signals. For the autoregressive baseline (Autoreg. 1/0), the results indicate that the model primarily relies on the autoregressive states while largely ignoring the less-informative measurements (see Figure 7). This provides strong indication of an over-reliance on the autoregressive components.
A related conclusion can be drawn by comparing the MSE of the two autoregressive baselines (Autoreg. 1/0 and 16/0) over rollout (see Figure 6), where both models exhibit very similar error trajectories. This suggests that the additional information provided by the longer measurement history in Autoreg 16/0 is not effectively used, indicating that the model overly relies on its autoregressive state.
Figure 7: Relative input sensitivity of the loss with respect to different input channel groups, measured using mean squared gradient norms with respect to the model input. Gradients are aggregated over spatial and batch dimensions and grouped by input type, as the conditioning for the autoregressive model (Autoreg. 1/0) is provided via the channel dimension. The gradients are normalized to obtain relative contributions for each conditioning signal, and the running mean (dashed line) and standard deviation (shaded area) of these quantities are visualized. Over the course of the training, sensitivity with respect to the measurement inputs decreases, while sensitivity to the autoregressive states increases. This could suggest, that the model increasingly relies on the more informative autoregressive context relative to the measurements.
@@
Stability through future predictions.
Comparing FlowPAINT 16/0 with the FlowPAINT 16/8 variant highlights the benefit of incorporating future-state predictions. This holds even if we use the same number of predicted frames per batch (see Appendix B.3 for more details).
Varying the number of measurement points.
We depict the MSE with varying probe measurements points in the grid setting in Figure 10. Although FlowPAINT was trained with 25 randomly resampled probe points for each datapoint, the best performance is obtained with around 100 probe points. The autoregressive UNet’s MSE stays approximately the same, even if the number of measurement points are increased.
4.4 Limitations
Sampling a sequence vs. a single state.
PAINT may be used for sampling sequences or individual timesteps. Sampling a connected, temporally smooth sequence, $\bm{x}_{[t-h,t]}\sim p(\bm{x}_{[t-h,t])}\mid\bm{m}_{[t-h,t]})$ aids result interpretation (see Figure 21). Sampling individual timesteps $\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{m}_{[t-h,t]})$, may be used for real-time state estimation in a sliding window context, albeit without the same temporal smoothness.
Discontinuities.
A fundamental limitation of parallel-in-time models is the inherently discontinuous temporal nature of the predicted trajectory. This lack of temporal smoothness makes generated trajectories less interpretable and complicates the analysis of temporal properties. However, this limitation is acceptable in scenarios where the primary objective is to obtain an estimate of the system’s current or near-future state. In our experiments, we also observe that for reasonably small time steps, the induced discontinuities relative to an autoregressive baseline are modest (see Figure 23 in the Appendix B.3).
@@ 4.4 Limitations
FlowPAINT needs substantially more training resources and is slower at inference than the autoregressive baseline. It was trained on 12 A100 (64 GB VRAM) for 30 hours, whereas the UNet was trained on a single H100 (80GB VRAM) for 24 hours. Notably, the UNet takes about $66$ milliseconds and FlowPAINT around $6.6$ seconds for 20 denoising steps. Further details can be found in Appendix B.2.
@@ 5 Conclusion
In this work, we introduced Parallel-in-time Neural Twins (PAINT), a novel data-driven framework designed to bridge the gap between static neural surrogates and adaptive digital replicas for real-time reconstruction of dynamical systems. PAINT employs a generative neural network to model the joint conditional probability of system states, enabling it to provably stay on-trajectory without reliance on initial conditions. One of our key findings is the problem of over-reliance on autoregressive state in standard autoregressive models. Our insights on this issue are supported by both theoretical analysis and experimental results.
The practical effectiveness of our approach was demonstrated through the application of FlowPAINT to 2D turbulent fluid dynamics, underscoring the potential of PAINT for complex, real-world systems. Taken together, this paper marks a significant advancement in the design of interpretable and widely applicable neural twins for intricate dynamical systems, paving the way for more accurate and adaptable modeling techniques.
@@
A major limitation of our method are the higher computational costs. Future work could explore more efficient architectures or alternative training paradigms, such as subsampling timesteps, to improve scalability. Regarding the observed over-reliance on the autoregressive state, we provide theoretical insights and empirically validate behavior in our experiments. However, these findings are specific to the considered data, task, and model and may not directly generalize to other settings. Another limitation is the sliding window approach, which prevents truly continuous generations. This could potentially be mitigated by stitching techniques (Wei et al., 2019).
@@ Acknowledgements
Appendix A Theory
A.1 Parallel-in-time models are on-trajectory
Here we provide a simple derivation to show that parallel-in-time models are on-trajectory under Assumption 3.1. We will show that for an appropriate window size $h$, for all $t\in\mathbb{N}$ and any $\epsilon_{1}>0$:
First, from Assumption 3.1 choose a window size $h$ such that:
Then, from the Universal Function Approximator Theorem (Cybenko, 1989), choose an approximation of the true distribution as follows:
A.2 In general, autoregressive models are not on-trajectory
do not generally fulfill the on-trajectory property. We construct this example while demonstrating the over-reliance on autoregressive state. Being on-trajectory is defined as: for all $t\in\mathbb{N}$ and all $\epsilon>0$:
From Section 3.3 it is known that autoregressive models accumulate errors with the product of Jacobians, similar to classical ODEs (Orrell, 2005).
Logistic map.
Let $f:\mathbb{R}\mapsto\mathbb{R}$ be the logistic map, which is chaotic for $r=3.8$ (May, 1976):
$f(x_{t-1},\eta_{t})$ $=r\cdot x_{t-1}\cdot(1-x_{t-1})$ (13)
Appendix B Experimental details
B.1 Dataset
@@
The dataset is generated using the incompressible, pressure-based solver pimpleFOAM in OpenFOAM 8 (OpenCFD (2009)), with subgrid scales modeled by the Smagorinsky LES model (Pope, 2001; Fröhlich, 2006). Simulations are performed on a structured mesh consisting of 19040 cells, and for training purposes, a subregion is extracted and interpolated onto a regular 128 × 128 grid to facilitate adaptability. The numerical simulations made use of a timestep size of $\Delta t=6.5\times 10^{-4}\text{s}$. For training purposes, the data are sampled every 70th timestep, resulting in an effective temporal resolution of $4.55\times 10^{-2}\text{s}$. For each Reynolds number, the dataset contains 1000 such snapshots.
@@
In reality, there is no perfectly two-dimensional turbulence; even quasi-2D flows, such as atmospheric flows, exhibit three-dimensional structures at smaller scales. Nevertheless, we are confident that the proposed method, having demonstrated robust performance in two-dimensional turbulence, will generalize effectively to three-dimensional turbulent flows. For a more detailed discussion of two-dimensional turbulence, see Davidson (2015).